# Depth-Wise Convolutions in Vision Transformers for Efficient Training on Small Datasets

### v3

@@ Abstract
The Vision Transformer (ViT) leverages the Transformer’s encoder to capture global information by dividing images into patches and achieves superior performance across various computer vision tasks. However, the self-attention mechanism of ViT captures the global context from the outset, overlooking the inherent relationships between neighboring pixels in images or videos. Transformers mainly focus on global information while ignoring the fine-grained local details. Consequently, ViT lacks inductive bias during image or video dataset training. In contrast, convolutional neural networks (CNNs), with their reliance on local filters, possess an inherent inductive bias, making them more efficient and quicker to converge than ViT with less data. In this paper, we present a lightweight Depth-Wise Convolution module as a shortcut in ViT models, bypassing entire Transformer blocks to ensure the models capture both local and global information with minimal overhead. Additionally, we introduce two architecture variants, allowing the Depth-Wise Convolution modules to be applied to multiple Transformer blocks for parameter savings, and incorporating independent parallel Depth-Wise Convolution modules with different kernels to enhance the acquisition of local information. The proposed approach significantly boosts the performance of ViT models on image classification, object detection and instance segmentation by a large margin, especially on small datasets, as evaluated on CIFAR-10, CIFAR-100, Tiny-ImageNet and ImageNet for image classification, and COCO for object detection and instance segmentation. The source code can be accessed at https://github.com/ZTX-100/Efficient_ViT_with_DW.
@@ 2.1 Vision Transformers
ViT [2] introduces the Transformer models into vision recognition by splitting the images into fixed-size patches and then tokenizing each patch into the token so that the image patches can be utilized in the attention module of Transformer models. Many variations and improvements have been proposed [4][5][6] and applied to various vision tasks such as point cloud completion [7] and crowd counting [8]. DeiT [9] employs distillation tokens for attention learning from the teacher models to the student models. CaiT [10] introduces LayerScale to effectively train the ViT models with deeper layers so that the performance of deep ViT models could be further boosted. Hierarchical Vision Transformer architecture [11][12][13][14] are designed to better suit vision tasks by reducing the size of feature maps as the network progresses deeper, resembling the structure of CNN architectures.
@@ 2.2 Vision Transformers and Convolutions
The computational structure of some previous models focuses on merging convolutional networks into the Multi-Head Self-Attention (MHSA) or Feed-Forward Network (FFN), making convolutional networks essential components of Transformer structures. Additionally, the convolutional parts in some of these works are not necessarily lightweight. In contrast to these previous studies, our approach aims to efficiently integrate Transformer blocks and convolutions with minimal computational overhead. It is designed as a versatile and straightforward module that can be easily incorporated into various Vision Transformer models.
@@ 3.2 Our Approach
Figure 2: The architecture variants of our proposed approach for bypassing multiple Transformer blocks. The upper, middle, and lower structures indicate that the Depth-Wise module bypasses 2, 3, and 4 Transformer blocks, respectively. For some Vision Transformer models with deeper layers, bypassing more blocks might be a good choice to reduce the parameters and computational costs.
@@ 3.3 Architecture Variants
Figure 3: The architecture variants of our method for multiple DWConv modules in parallel. Multiple independent parallel DWConv modules operate on Transformer blocks concurrently to capture local details simultaneously. This structure could be combined with previous variants shown in Fig. 2 to include N Transformer blocks in the DWConv modules.
Furthermore, the DWConv modules could operate in parallel with the same or different kernels to capture the local information independently, as demonstrated in Fig. 3. In the experiments, we leverage parallel DWConv modules with different kernel sizes to illustrate the performance of the variants. To further reduce the number of parameters and computational cost, multiple independent parallel DWConv modules could be combined with the aforementioned variants so that multiple Transformer blocks are contained by the DWConv modules. In Fig. 3, $N$ Transformer blocks are encompassed in the DWConv modules and $N\geq 1$.
Not modifying the structures of MHSA and FFN makes our approach more flexible for use with most Transformer models, rather than being designed for specific ones. The proposed architecture variants illustrate the flexibility of our methods compared to some existing hybrid architectures that combine convolutions and Transformers. The Transformers are greatly enhanced by the proposed DWConv module with minimal overhead. Additionally, the structure can be easily modified to further enhance performance or save parameters and computations with different variants.
@@ 4.1 Classification Performance on Small Datasets
CIFAR10 [38], CIFAR100 [38], and Tiny-ImageNet [39] are exploited as small datasets for training and evaluating image classification tasks.
@@ 4.1.3 CaiT
Figure 4: Some Grad-CAM visualization with ViT-Tiny and Swin-Tiny models. The vanilla Transformer models tend to capture the global information, which could be illustrated in the CAM visualization. With our method, the models could capture both local details and global views, especially when the objects are relatively small.
@@ 4.2 Classification Performance on ImageNet-1K
Model Accuracy Params FLOPs
ResNet-18 [52] 71.5 11.7M 1.8G
ResNet-34 [52] 76.4 21.8M 3.7G
ResNet-50 [52] 80.4 25.6M 4.1G
SE-ResNet-50 [53] 80.0 28.1M 4.1G
DeiT-Ti [9][26] 72.2 5.7M 1.3G
Visformer-Ti [54] 78.6 10.3M 1.3G
PVT-Tiny [11] 75.1 13.2M 1.9G
DeiT-S [9][26] 79.9 22.1M 4.6G
PVT-Small [11] 79.8 24.5M 3.8G
MSG-T [16] 80.9 28M 4.6G
DiT-B1 [55] 79.9 30.3M 2.0G
Visformer-S [54] 81.5 40.2M 4.9G
CaiT-xxs12 74.85 6.6M 1.3G
CaiT-xxs24 77.66 11.9M 2.5G
Swin-Tiny 81.14 28.3M 4.5G
CaiT-xxs12 (ours) 75.89 (+1.04) 6.6M 1.3G
CaiT-xxs24 (ours) 79.66 (+2.00) 12.0M 2.5G
Swin-Tiny (ours) 81.73 (+0.59) 28.3M 4.5G
We utilize CaiT [10] and Swin Transformer [15] to illustrate the performance of our approach on ImageNet. The kernel size for our method is $3\times 3$ and the DWConv module is applied to each Transformer block. When our approach is applied to the models, the positional embeddings and talking heads attention in CaiT are retained for better accuracy, while LayerScale is eliminated. For Swin Transformer, our approach is directly applied to the model without any other changes. The experimental results are demonstrated in Table 6. We employ top 1 accuracy to measure the performance of the models and the results of ResNet models are from [56]. The performance of CaiT and Swin-Transformer on ImageNet-1K is further boosted (up to 2%) by our method.
@@ 4.3 Object Detection and Instance Segmentation
For the experimental settings, we employ AdamW [43] as the optimizer and the warmup iterations are 500. We utilize four NVIDIA P100 GPUs to train the model with 2 samples each GPU. The initial learning rate is set at 5e-5 and the learning rate is reduced by 10 at epochs 9 and 12, respectively. The image scale for the experiments is 1333$\times$800. The total epochs for the experiments are 12.

### v4

@@ Abstract
The Vision Transformer (ViT) leverages the Transformer’s encoder to capture global information by dividing images into patches and achieves superior performance across various computer vision tasks. However, the self-attention mechanism of ViT captures the global context from the outset, overlooking the inherent relationships between neighboring pixels in images or videos. Transformers mainly focus on global information while ignoring the fine-grained local details. Consequently, ViT lacks inductive bias during image or video dataset training. In contrast, convolutional neural networks (CNNs), with their reliance on local filters, possess an inherent inductive bias, making them more efficient and quicker to converge than ViT with less data. In this paper, we present a lightweight Depth-Wise Convolution module as a shortcut in ViT models, bypassing entire Transformer blocks to ensure the models capture both local and global information with minimal overhead. Additionally, we introduce two architecture variants, allowing the Depth-Wise Convolution modules to be applied to multiple Transformer blocks for parameter savings, and incorporating independent parallel Depth-Wise Convolution modules with different kernels to enhance the acquisition of local information. The proposed approach significantly boosts the performance of ViT models on image classification, object detection, and instance segmentation by a large margin, especially on small datasets, as evaluated on CIFAR-10, CIFAR-100, Tiny-ImageNet and ImageNet for image classification, and COCO for object detection and instance segmentation. The source code can be accessed at https://github.com/ZTX-100/Efficient_ViT_with_DW.
@@ 2.1 Vision Transformers
ViT [2] introduces the Transformer models into vision recognition by splitting the images into fixed-size patches and then tokenizing each patch into the token so that the image patches can be utilized in the attention module of Transformer models. Many variations and improvements have been proposed [4][5][6] and applied to various vision tasks such as point cloud completion [7] and crowd counting [8]. DeiT [9] employs distillation tokens for attention learning from the teacher models to the student models. CaiT [10] introduces LayerScale to effectively train the ViT models with deeper layers so that the performance of deep ViT models could be further boosted. Hierarchical Vision Transformer architecture [11][12][13][14][15] are designed to better suit vision tasks by reducing the size of feature maps as the network progresses deeper, resembling the structure of CNN architectures.
@@ 2.2 Vision Transformers and Convolutions
The computational structures of some previous models focus on integrating convolutional networks into the Multi-Head Self-Attention (MHSA) or Feed-Forward Network (FFN), making convolutional networks essential components of Transformer architectures. Additionally, the convolutional components in some of these studies are not necessarily lightweight. In contrast, our approach aims to efficiently combine Transformer blocks with convolutions while minimizing computational overhead. It is designed as a versatile and straightforward module that can be easily integrated into various Vision Transformer models.
@@ 3.2 Our Approach
Figure 2: The architecture variants of our proposed approach involve bypassing multiple Transformer blocks. Structures (a), (b), and (c) represent the Depth-Wise module bypassing 2, 3, and 4 Transformer blocks, respectively. For Vision Transformer models with deeper layers, bypassing additional blocks may be a beneficial strategy to reduce both parameters and computational costs.
@@ 3.3 Architecture Variants
Figure 3: The architecture variants of our method involve multiple DWConv modules operating in parallel. These independent DWConv modules, each with different kernel sizes, run concurrently on Transformer blocks to capture local details simultaneously. This structure can be combined with previous variants shown in Fig. 2 to include N Transformer blocks in the DWConv modules.
Furthermore, the DWConv modules could operate in parallel with various kernel sizes to capture the local information independently, as illustrated in Fig. 3. In the experiments, we leverage parallel DWConv modules with different kernel sizes to demonstrate the performance of the variants. To further reduce the number of parameters and computational cost, multiple independent parallel DWConv modules could be combined with the aforementioned variants so that multiple Transformer blocks are contained by the DWConv modules. In Fig. 3, $N$ Transformer blocks are encompassed in the DWConv modules and $N\geq 1$.
Not modifying the structures of MHSA and FFN makes our approach more flexible for use with most Transformer models, rather than being designed for specific ones. The proposed architecture variants illustrate the flexibility of our methods compared to some existing hybrid architectures that combine convolutions and Transformers. The Transformers are greatly enhanced by the proposed DWConv module with minimal overhead. Additionally, the structure can be easily modified to further enhance the performance or save the parameters and computations with different variants.
@@ 4.1 Classification Performance on Small Datasets
CIFAR10 [39], CIFAR100 [39], and Tiny-ImageNet [40] are exploited as small datasets for training and evaluating image classification tasks. The classification accuracy is defined as the ratio of correctly classified samples to the total number of samples. In our paper, we use Top-1 accuracy for classification.
@@ 4.1.3 CaiT
Figure 4: Some Grad-CAM visualization with ViT-Tiny and Swin-Tiny models. The vanilla Transformer models tend to capture the global information, as illustrated in the CAM visualizations. With our method, the models are able to capture both local details and global perspectives, particularly when dealing with smaller objects. Please note that the original images in the figure are from the Tiny-ImageNet dataset with a low resolution of 64$\times$64 pixels. Thus, they appear blurred when enlarged.
@@ 4.2 Classification Performance on ImageNet-1K
Model Accuracy Params FLOPs Year
ResNet-18 [53][54] 71.5 11.7M 1.8G 2016
ResNet-34 [53][54] 76.4 21.8M 3.7G 2016
ResNet-50 [53][54] 80.4 25.6M 4.1G 2016
SE-ResNet-50 [55][54] 80.0 28.1M 4.1G 2018
DeiT-Ti [9][27] 72.2 5.7M 1.3G 2021
Visformer-Ti [56] 78.6 10.3M 1.3G 2021
PVT-Tiny [11] 75.1 13.2M 1.9G 2021
DeiT-S [9][27] 79.9 22.1M 4.6G 2021
PVT-Small [11] 79.8 24.5M 3.8G 2021
MSG-T [17] 80.9 28M 4.6G 2022
DiT-B1 [57] 79.9 30.3M 2.0G 2023
Visformer-S [56] 81.5 40.2M 4.9G 2021
CaiT-xxs12 74.85 6.6M 1.3G -
CaiT-xxs24 77.66 11.9M 2.5G -
Swin-Tiny 81.14 28.3M 4.5G -
CaiT-xxs12 (ours) 75.89 (+1.04) 6.6M 1.3G -
CaiT-xxs24 (ours) 79.66 (+2.00) 12.0M 2.5G -
Swin-Tiny (ours) 81.73 (+0.59) 28.3M 4.5G -
We utilize CaiT [10] and Swin Transformer [16] to illustrate the performance of our approach on ImageNet. The kernel size for our method is $3\times 3$ and the DWConv module is applied to each Transformer block. When our approach is applied to the models, the positional embeddings and talking heads attention in CaiT are retained for better accuracy, while LayerScale is eliminated. For Swin Transformer, our approach is directly applied to the model without any other changes. The experimental results are demonstrated in Table 6. We employ top-1 accuracy to measure the performance of the models and the results of ResNet models are extracted from [54]. The performance of CaiT and Swin-Transformer on ImageNet-1K is further boosted (up to 2%) by our method.
@@ 4.3 Object Detection and Instance Segmentation
For the experimental settings, we employ AdamW [44] as the optimizer and the warmup iterations are 500. We utilize four NVIDIA P100 GPUs to train the model with 2 samples for each GPU. The initial learning rate is set at 5e-5 and the learning rate is reduced by 10 at epochs 9 and 12, respectively. The image scale for the experiments is 1333$\times$800. The total epochs for the experiments are 12.